1) In a population genetic screening study for an autosomal recessive disease, you observe an allelic frequency of 0.8 for the dominant allele and 0.2 for the disease-causing recessive allele. What is the frequency of the disease (phenotypic) in the population? show work (i.e. Punnett's squares or other) especially
Answer #1

According to Hardy Weinberg equilibrium p2 + q2 + 2pq =1 where p is the frequency of the dominant allele and q is the frequency of the recessive allele.

p2  is the frequency of homozygous dominant individuals , q2 is the frequency of homozygous recessive individuals and pq is the frequency of heterozygous individuals.

If a disease has an autosomal recessive inheritance pattern, it means that only individuals with two recessive alleles, i.e individuals with a homozygous recessive genotype are affected. Individuals that are homozygous dominant are not diseased and those that are heterozygous are unaffected carriers.

Therefore, the frequency of the disease in the population = q2 = 0.2*0.2 =0.04. That is 4%(0.04*100) of the population is diseased.